Suppose that there are two goods, X and Y . Thomasíutility function is
U = XY + X. The price of X is P, and the price of Y is $40. Income is
$200. Derive his individual demand for X as a function of P.
1
Expert's answer
2013-10-02T11:35:30-0400
Suppose that there are two goods, X and Y. Thomas utility function is
U = XY + X. The price of X is P, and the price of Y is $40. Income is
$200. Derive his individual demand for X as a function of P.

Income function is: I = Px*X + Py*Y = 200
P*X + 40Y = 200
The individual demand for X as a function of P will be:
Qd = (200 - 40Y)/P